# EraseDiff: Erasing Data Influence in Diffusion Models

### 11 Jan 2024 v1

@@ 5.1 Setup
Unscrubbed 19.62 12.05 0.8622 8.87 0.7614
EraseDiff (Ours) 256.27 ($+236.65$) 294.08 ($+282.03$) 0.0026 330.33 ($+321.46$) 0.0000
@@
5.2 Effectiveness of data removal
We aim to unlearn images conditioned on classes of birds and ships (represented by $c=2$ and $c=8$) on CIFAR10, faces with the last ethnicity labeled as Indian ($c=3$) on UTKFace, and images with attribute blond hair (unconditional case) on CelebA and CelebA-HQ, respectively. Results in different scenarios can be found in Appendix A.2. For an effective unlearning algorithm, the scrubbed models are expected to contain little information about the forgetting data. In this section, we first show the effectiveness of our proposed unlearning algorithm on diffusion models.
FID score. Tab. 1 presents the results of FID on the generated images conditioned on the forgetting classes ${\mathcal{C}}_{f}$. FID scores are computed between the generated images and the corresponding images with the same labels from the training data ${\mathcal{D}}_{f}$. As shown in Tab. 1, generated images conditioned on ${\mathcal{C}}_{f}$ from the models scrubbed by our proposed unlearning algorithm have larger FID scores than those from the unscrubbed models, indicating that our algorithm successfully erases the information of the forgetting data carried by the models. For example, on CIFAR10, the FID score of generated images conditioned on the label birds increased from 19.62 to 256.27, indicating that these generated bird images are notably dissimilar from the bird images present in the training data.
Accuracy. The classification accuracy of the generated images conditioned on ${\mathcal{D}}_{f}$ decreased from around 0.86 to 0.002 on CIFAR10, and from 0.76 to 0 on UTKFace, respectively. This demonstrates that, among the generated 10K images on CIFAR10 and 25K images on UTKFace, only 26 images are classified into the categories of birds or ships on CIFAR10 and no faces are classified as Indian celebrities, further verifying the effectiveness of our unlearning algorithm.
Loss distribution. To further measure the effectiveness of data removal, we report the results of MIA accuracy to indicate whether the data is used to train the models. The attack model is a binary classifier and is trained using the loss value w.r.t. the forgetting data, and data from the test set. As shown in Fig. 1, after scrubbing the models using our proposed unlearning algorithm EraseDiff, the distribution of loss values on ${\mathcal{D}}_{f}$ is similar to that on the test data, indicating that our methodology successfully erased the information of the forgetting data carried by the model.
Ideally, the loss values w.r.t. the training data would be lower than those w.r.t. the data from the test set. However, diffusion models have impressive capacities to generate diverse and high-quality images, leading to difficulty in distinguishing between the generated images and the real images in the training data. For example, on CIFAR10 shown in Fig. 1, the distribution of loss values on ${\mathcal{D}}_{f}$ is quite similar to that on the unseen data for the unscrubbed model, resulting in the instability of MIA performance. Advanced methods to measure MIA could be leveraged, we leave it for future studies.
Figure 2: Images generated by conditional DDIM. Images in the green dashed box are generated by conditioning on the remaining labels ${\mathcal{C}}_{r}$ and those in the red solid box are generated by conditioning on the forgetting classes ${\mathcal{C}}_{f}$. Our unlearning algorithm could successfully scrub the information of ${\mathcal{C}}_{f}$ carried by the models while maintaining the model utility over ${\mathcal{C}}_{r}$.
Figure 3: KL distance between the approximator output ${\bm{\epsilon}}_{T}$ and standard Gaussian noise ${\bm{\epsilon}}$. Our unlearning algorithm maintains the model utility on ${\mathcal{D}}_{r}$ while ensuring that the predicted distribution significantly diverges from the predefined standard Gaussian noise distribution for ${\mathcal{D}}_{f}$.
KL divergence. Here, we propose computing the KL divergence distance between the model output distribution and the pre-defined Gaussian noise distribution. The diffusion process would add noise to the clean input ${\mathbf{x}}$, consequently, the model output ${\bm{\epsilon}}_{T}$ distribution given these noisy images ${\mathbf{x}}_{t}$ tends to converge towards the Gaussian noise distribution ${\bm{\epsilon}}$. As depicted in Fig. 3, the y-axis represents the number of occurrences of the specific distance value. We can see that the distance between the output distribution w.r.t. ${\mathcal{C}}_{r}$ and the Gaussian distribution for our scrubbed model is close to that of the unscrubbed model, while the distance between the output distribution w.r.t. the ${\mathcal{C}}_{f}$ is further away from the Gaussian distribution for our scrubbed model than the unscrubbed model. This means for ${\mathcal{D}}_{r}$, our scrubbed models exhibit behavior akin to the unscrubbed models, while for ${\mathcal{D}}_{f}$, our scrubbed models would generate images significantly deviate from ${\bm{\epsilon}}$. These results underscore the efficacy of our unlearning algorithm in scrubbing the unscrubbed model concerning the information of ${\mathcal{D}}_{f}$, all the while preserving the model utility for ${\mathcal{D}}_{r}$.
Visualizations. To provide a more tangible assessment of unlearning effectiveness, we visualize some generated images. Fig. 2 showcases samples generated by the conditional DDIM models from Tab. 1. When conditioned on ${\mathcal{D}}_{f}$, the scrubbed models predominantly produce unrecognizable images. More examples and examples of the unconditional case can be found in Appendix A.2.
Table 2: FID score over the remaining classes on conditional DDIM. FID score is evaluated on 50K generated images (5K per class for CIFAR10, 12.5K per class for UTKFace). The quality of the generated image conditioned on the remaining classes drops a little after scrubbing (examples in Fig. 2) but it is acceptable given the significant unlearning impact of the forgetting classes.
CIFAR10 UTKFace $c=0\downarrow$ $c=1\downarrow$ $c=3\downarrow$ $c=4\downarrow$ $c=5\downarrow$ $c=6\downarrow$ $c=7\downarrow$ $c=9\downarrow$ $c=0\downarrow$ $c=1\downarrow$ $c=2\downarrow$ Unscrubbed 17.04 9.67 19.88 14.78 20.56 17.16 11.53 11.44 7.37 11.28 9.72 EraseDiff (Ours) 29.61 22.10 28.65 27.68 35.59 23.93 21.24 24.85 8.08 13.52 12.37
5.3 Efficacy of preserving model utility.
For an effective unlearning algorithm, the scrubbed models are also expected to maintain utility over the remaining classes. So we present the FID scores over the generated images conditioned on ${\mathcal{C}}_{r}$ from the scrubbed models in Tab. 2. The FID scores increase compared with the generated images from the original models. As shown in Fig. 2, it’s evident that the quality of generated images experiences a slight decrease. However, given the substantial unlearning impact of the forgetting classes, this drop in quality is acceptable. Furthermore, as shown in Figs. 1 and 3, our scrubbed models could obtain similar results as the unscrubbed models over the remaining classes.
Table 3: Results on CIFAR10 trained with conditional DDIM. Although NegGrad and BlindSpot could scrub the information of the forgetting classes, they cannot maintain the model utility. Other methods fail to scrub the relevant information completely. Examples are shown in Fig. 4.

### 05 Feb 2024 v2

@@ 5.1 Setup
5.2 Results on DDPM and DDIM
Following SA, we aim to forget the ‘airplane’ class on CIFAR10 with conditional DDPM, we use the checkpoint provided in SA. We further conduct experiments when we have no access to ${\mathcal{D}}_{r}$. Besides, we aim to unlearn images conditioned on classes of birds and ships (represented by $c=2$ and $c=8$) on CIFAR10 with conditional DDIM.
DDPM results. Tabs. 1 and 2 present results when forgetting the ‘airplane’ class. FID scores are computed between the generated 45K images and the corresponding ground truth images with the same labels from ${\mathcal{D}}_{r}$. The FID scores decrease slightly compared with the generated images from the original models by $\sim$ 1.7; the quality of the generated images experiences a slight improvement. However, there is a decrease in recall (diversity), which can be attributed to our scrubbed model being fine-tuned over ${\mathcal{D}}_{r}$, suggesting a tendency towards overfitting. The probability of classifying the images of the ‘airplane’ class drops around 75%, and the entropy increases significantly, indicating that more information in the generated images conditioned on the ‘airplane’ class has been erased.
No access to ${\mathcal{D}}_{r}$. As presented in SA Heng & Soh (2023b), we further consider having no access to the remaining data and apply GR to generate 4.5K samples to be ${\mathcal{D}}_{r}^{{}^{\prime}}$. As illustrated in Tab. 1, while SA outperforms EraseDiff in terms of the quality of generated images, it’s important to note the difference in training duration: our model was trained for only 200 steps, in contrast to SA’s extensive training over 20K steps. We hypothesize that the observed drop in quality can be mitigated by increasing the GR sample size and enhancing the quality of the generated images. However, we will explore and validate this hypothesis in future investigations.
Overhead. Assuming that the computational complexity over the data ${\mathcal{D}}$ is represented as ${\mathcal{O}}(Nl(n))$, similarly, that over the remaining data ${\mathcal{D}}_{r}$ is denoted as ${\mathcal{O}}(N_{r}l(n))$. The computational complexity of SA involves two distinct stages: the computation of FIM and the subsequent forgetting stage. We consider the maximum memory usage across both stages, the metric ‘Time’ is exclusively associated with the duration of the forgetting stage for SA. Tab. 2 show that EraseDiff outperforms SA in terms of efficiency, achieving a speed increase of $>2$ times for a single training step. This is noteworthy, especially considering the necessity for computing FIM in SA for different datasets and models.
DDIM results. Next, we conduct experiments with DDIM when forgetting the ‘bird’ and ‘ship’ classes. Tabs. 3 and 3 present the results. FID scores are computed for 5K generated images conditioned on each class and the number of time steps for sampling is set to be 100. As illustrated in Tab. 3, the NegGrad model and BlindSpot model can erase the information related to forgetting classes ${\mathcal{C}}_{f}$, but struggle to retain the model utility. For example, images generated conditioned on ${\mathcal{C}}_{f}$ (birds and ships) yield the FID score exceeding 200, and $\sim$2000 of these generated images are classified into the birds or ships categories. However, for the generated images conditioned on the remaining classes ${\mathcal{C}}_{r}$, the FID scores also surpass 200, as visualized in Fig. 3, these generated images appear corrupted. Other baseline models can preserve the model utility over ${\mathcal{C}}_{r}$ but fall short in sufficiently erasing the information regarding ${\mathcal{C}}_{f}$. Our proposed unlearning algorithm, instead, adeptly scrubs pertinent information while upholding model utility. Furthermore, we can observe that our scrubbed model has a WD of 1.3534, while the minimal WD is 1.3533, indicating that our scrubbed models are in close alignment with the retrained model where the forgetting data ${\mathcal{D}}_{f}$ never attends in the training process.
5.3 Results on Stable Diffusion
[…]
Table 4: Results on Imagenette with Stable Diffusion.
SD v1.4 ESD-x EraseDiff-x
FID $\downarrow$ 4.8883 3.0824 3.0902
$P_{\psi}({\mathbf{y}}=c_{f}|{\mathbf{x}}_{f})$$\downarrow$ 0.74 0.00 0.00
Forget class. When performing class-wise forgetting, we fine-tune the cross-attention layers in SD. Following Fan et al. (2023), we set the prompt as ‘an image of [$c$]’. For the forgetting class ‘tench’, we choose the ground truth backward distribution to be a class other than ‘tench’, denoted as EraseDiff-x. We generate 100 images for each prompt. As shown in Tab. 4, when the probability of classifying the generated images of the ‘tench’ class is 0%, the FID score would increase to 3.0902, slightly larger (ie., $<0.008$) than ESD-x. Results of EraseDiff reported in Tab. 4 is obtained over 58 iterations. When updated for 78 iterations, the probability of classifying the generated images of the ‘tench’ class is 5%, but EraseDiff-x achieves the lowest FID score of the generated images conditioned on $c_{r}$, which is 2.5932, dropped by around 0.5 compared to ESD-x. When we choose the uniform distribution, the probability of classifying the generated images of the ‘tench’ class is 0%, but the FID score increases up to $\sim$3.87, indicating that setting the ground truth backward distribution to be a uniform distribution could affect the remaining classes for the large-scale datasets. Note that in Tab. 4, ESD-x and EraseDiff-x outperform the original model SD v1.4. This could be due to the former methods performing fine-tuning.
Figure 4: Quantity of nudity content detected using the NudeNet classifier from I2P data and 1K sampled images. Our method effectively erases nudity content from Stable Diffusion, outperforming SD v2.1 and ESD-u.
Figure 5: Generated images with the prompt “a photo of a naked person” when forgetting the concept of nudity.
Forget nudity. Next, we attempt to forget the concept of nudity in SD v1.4. We only fine-tune the unconditional layers in SD v1.4, denoted as EraseDiff-u. For all models, 4703 images are generated using I2P prompts, and 1K images are generated using the prompts {‘nudity’, ‘naked’, ‘erotic’, ‘sexual’} Heng & Soh (2023b), We set the ground truth backward distribution with the prompt “a person wearing clothes”. The quantity of nudity content is detected using the NudeNet classifier. Figs. 4 and 5 present the results for SD v2.1 (trained on a dataset filtered for nudity), ESD-u, and EraseDiff-u. In Fig. 4, the number in the y-axis denotes the number of exposed body parts generated by the SD v1.4 model. Fig. 4 presents the percentage change in exposed body parts w.r.t. SD v1.4. We can find that, EraseDiff-u outperforms other methods, it reduces the amount of nudity content compared to SD v1.4, SD v2.1, and ESD-u, particularly on sensitive content like Female/Male Breasts and Female/Male Genitalia.
6 Conclusion and Limitations
In this work, we explored the unlearning problem in diffusion models and proposed an efficient unlearning method EraseDiff. Comprehensive experiments on unconditional and conditional diffusion models demonstrate the proposed algorithm’s effectiveness in data removal, its efficacy in preserving the model utility, and its efficiency in unlearning. We hope the proposed approach could serve as an inspiration for future research in the field of diffusion unlearning.
The scrubbed models could be biased for generation, but we do not take this into account in the experiments. Future directions for diffusion unlearning could include assessing fairness post-unlearning and using advanced privacy-preserving training techniques. Besides, when having no access to the training data, with generated images for unlearning, the image quality after scrubbing the model using EraseDiff would be worse; further work could investigate methods to improve generated image quality without access to the training data.
Impact Statements
DMs have experienced rapid advancements and have shown the merits of generating high-quality data. However, concerns have arisen due to their ability to memorize training data and generate inappropriate content, thereby negatively affecting the user experience and society as a whole. Machine unlearning emerges as a valuable tool for correcting the algorithms and enhancing user trust in the respective platforms. It demonstrates a commitment to responsible AI and the welfare of its user base. However, while unlearning protects privacy, it may also hinder the ability of relevant systems, potentially lead to biased outcomes, and even be adopted for malicious usage.
Appendix A Implementation Details
Algorithm 2 EraseDiff.
0:  Well-trained model $\epsilon_{{\bm{\theta}}_{0}}$, forgetting data ${\mathcal{D}}_{f}$ and subset of remaining data ${\mathcal{D}}_{rs}\subset{\mathcal{D}}_{r}$, outer iteration number $E$ and inner iteration number $K$, learning rate $\eta$ and hyparameter $\lambda$.
0:  Parameters ${\bm{\theta}}^{*}$ for the scrubbed model.
1:  for iteration s in E do
2:     $\bm{\phi}_{s}^{0}={\bm{\theta}}$.
3:     Get $\bm{\phi}_{s}^{K}$ by $K$ steps of gradient descent on $f(\bm{\phi}_{s},{\mathcal{D}}_{f})$ start from $\bm{\phi}_{s}^{0}$:
4:                  Sample $\{{\mathbf{x}}_{0},c\}\subset{\mathcal{D}}_{f},t\sim\text{Uniform}({1,\cdots,T}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$,
5:                  Compute $\nabla_{\bm{\phi}_{s}^{k}}\|\hat{{\bm{\epsilon}}}-\epsilon_{{{\bm{\phi}_{s}^{k}}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\hat{{\bm{\epsilon}}},t,c)\|$.
6:                  Get the constant loss ${\mathcal{L}}_{cs}=\|\hat{{\bm{\epsilon}}}-\epsilon_{{{\bm{\phi}_{s}^{K}}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\hat{{\bm{\epsilon}}},t,c)\|$ if $k=K$.
7:     Set the approximation:
8:                  Sample $\{{\mathbf{x}}_{0},c\}\subset{\mathcal{D}}_{f},t\sim\text{Uniform}({1,\cdots,T}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$,
9:                  Compute the loss ${\mathcal{L}}_{f}=\|\hat{{\bm{\epsilon}}}-\epsilon_{{\bm{\theta}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\hat{{\bm{\epsilon}}},t,c)\|-{\mathcal{L}}_{cs}$.
10:     Update the model:
11:                  Sample $\{{\mathbf{x}}_{0},c\}\subset{\mathcal{D}}_{rs},t\sim\text{Uniform}({1,\cdots,T}),{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$,
12:                  Compute the loss ${\mathcal{L}}_{s}=\|{\bm{\epsilon}}-\epsilon_{{{{\bm{\theta}}}}}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t,c)\|+\lambda{\mathcal{L}}_{f}$,
13:                  Update ${\bm{\theta}}_{s+1}={\bm{\theta}}-\eta\nabla_{{\bm{\theta}}}{\mathcal{L}}_{s}$.
14:  end for
Algorithm 3 BlindSpot Unlearning (Tarun et al., 2023b).
0:  A well-trained model $\epsilon$ with parameters ${\bm{\theta}}_{0}$, a randomly initialized blind model $\epsilon_{\bm{\psi}}(\cdot)$, forgetting data ${\mathcal{D}}_{f}$, remaining data ${\mathcal{D}}_{r}$ and all training data ${\mathcal{D}}={\mathcal{D}}_{f}\cup{\mathcal{D}}_{r}$. The learning rate $\zeta$, the number of epochs $E_{r}$ and $E_{u}$, hyper-parameter $\lambda$.
0:  Parameters ${\bm{\theta}}^{*}$ for the scrubbed model.
1:  Initialization ${\bm{\theta}}={\bm{\theta}}_{0}$.
2:  for $1,2,\dots,E_{r}$ do
3:     train the blind model $\epsilon_{\bm{\psi}}(\cdot)$ with the remaining data ${\mathcal{D}}_{r}$.
4:  end for
5:  for $1,2,\dots,E_{u}$ do
6:     for $({\mathbf{x}}_{i},c_{i})\in{\mathcal{D}}$ do
7:        $l^{i}_{f}=1$ if $({\mathbf{x}}_{i},c_{i})\in{\mathcal{D}}_{f}$ else $l^{i}_{f}=0$.
8:        ${\bm{\epsilon}}_{t}=\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{i},t,c_{i})$, where $t$ is the timestep and $t\in[1,T]$.
9:        ${\mathcal{L}}_{r}={\mathcal{L}}({\bm{\epsilon}}_{t},{\bm{\epsilon}})$ and ${\mathcal{L}}_{f}={\mathcal{L}}({\bm{\epsilon}}_{t},\epsilon_{\bm{\psi}}({\mathbf{x}}_{i},t,c_{i}))$.
10:        ${\mathcal{L}}_{a}=\lambda\sum_{j=1}^{k}\|\text{act}_{j}^{{\bm{\theta}}}-\text{act}_{j}^{\bm{\psi}}\|$, where $\text{act}_{j}$ is the output of each block in the UNet.
11:        ${\mathcal{L}}=(1-l^{i}_{f}){\mathcal{L}}_{r}+l^{i}_{f}({\mathcal{L}}_{f}+{\mathcal{L}}_{a})$.
12:        ${\bm{\theta}}={\bm{\theta}}-\zeta\frac{\partial{\mathcal{L}}}{\partial{\bm{\theta}}}$.
13:     end for
14:  end for
DDPM and DDIM.
Results on conditional DDPM follow the setting in SA Heng & Soh (2023b). We adopt the pre-trained DDPM from SA Heng & Soh (2023b). The batch size is set to be 128, the learning rate is $1\times 10^{-4}$, our model is trained for 200 training steps, and $N_{rs}=|{\mathcal{D}}_{rs}|=4.5K$. 5K images per class are generated for evaluation. For the remaining experiments, four and five feature map resolutions are adopted for CIFAR10 where image resolution is $32\times 32$, UTKFace, and CelebA where image resolution is scaled to $64\times 64$, respectively. The well-trained unconditional DDPM models on CIFAR10222https://huggingface.co/google/ddpm-cifar10-32 and CelebA-HQ [^3] https://huggingface.co/google/ddpm-ema-celebahq-256 are downloaded from Hugging Face. We used A40 and A100 for all experiments. All models apply the linear schedule for the diffusion process. We set the batch size $B=512$, $B=128$, $B=64$, $B=16$ for CIFAR10, UTKFace and CelebA, CelebA-HQ respectively. The linear schedule is set from $\beta_{1}=10^{-4}$ to $\beta_{T}=0.02$, the inference time step for DDIM is set to be 100, the guidance scale $w=0.1$, and the probability $p_{uncond}=0.1$ for all models. For the Unscrubbed and Retrain models, the learning rate is $3\times 10^{-4}$ for CIFAR10 and $2\times 10^{-4}$ for other datasets. We train the CIFAR10 model for 2000 epochs, the UTKFace and CelebA models for 500 epochs. For Finetune models, the learning rate is $3\times 10^{-4}$ for CIFAR10 and $2\times 10^{-4}$ for other datasets, all the models are finetuned on the remaining data ${\mathcal{D}}_{r}$ for 100 epochs. For NegGrad models, the learning rate is $1\times 10^{-6}$ and all the models are trained on the forgetting data ${\mathcal{D}}_{f}$ for 5 epochs. For BlindSpot models, the learning rate is $2\times 10^{-4}$. The partially-trained model is trained for 100 epochs on the remaining data ${\mathcal{D}}_{r}$ and then the scrubbed model is trained for 100 epochs on the data ${\mathcal{D}}$. For our scrubbed models, $N_{rs}=|{\mathcal{D}}_{rs}|\approx 8K$, the learning rate is $1\times 10^{-6}$ CelebA-HQ and $2\times 10^{-4}$ for other datasets.
SD.
We use the open-source SD v1.4 checkpoint as the pre-trained model for all SD experiments. The learning rate is $1\times 10^{-5}$, and we tune only the unconditional (non-cross-attention) layers of the latent diffusion model when erasing the concept of nudity, and conditional layers when forgetting the class in Imagenette. Our method is trained for 200 epochs. When forgetting nudity, we generate around 400 images with the prompts {‘nudity’, ‘naked’, ‘erotic’, ‘sexual’} and around 100 images with the prompt ‘a person wearing clothes’ to be the training data. We evaluate over 1K generated images for the Imagenette and Nude datasets. 4703 generated images with I2P prompts are evaluated using the open-source NudeNet classifier Bedapudi (2019).
Appendix B More Results
In the following, we present the results of Ablation studies, results when replacing ${\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$ with $\hat{{\bm{\epsilon}}}_{t}\sim{\mathcal{N}}(\bm{0.5},\mathbf{I}_{d})$ for LABEL:eq:obj_df_l, results when sampling from the uniform distribution $\hat{{\bm{\epsilon}}}_{t}\sim{\mathcal{U}}({\bm{0}},{\bm{1}})$, and results when trying to erase different classes/races/attributes under the conditional and unconditional scenarios. In brief, with more remaining data during the unlearning process, the generated image quality over the remaining classes ${\mathcal{C}}_{r}$ would be better while those over the forgetting classes ${\mathcal{C}}_{f}$ would be worse.
Table 5: Quantity of nudity content detected using the NudeNet classifier on 1K images generated with the prompts {‘nudity’, ‘naked’, ‘erotic’, ‘sexual’}.
Buttocks Female breast Female genitalia Male breast Male genitalia Feet Armpits Belly Anus SD v1.4 92 1496 351 68 89 792 686 58 0 SD v2.1 79 830 85 209 283 532 668 1 0 ESD-u 1 5 0 0 0 4 2 0 0 EraseDiff-u 0 0 0 0 1 0 1 0 0
Table 6: Quantity of nudity content detected using the NudeNet classifier on the I2P benchmark dataset (4703 images).
Buttocks Female breast Female genitalia Male breast Male genitalia Feet Armpits Belly Anus SD v1.4 28 276 13 44 55 158 162 7 0 SD v2.1 19 177 13 58 64 105 164 2 0 ESD-u 7 21 2 11 22 29 23 5 0 EraseDiff-u 2 7 0 0 7 4 8 0 0
Figure 6: Generated images with the prompt “nude” when forgetting the concept of nudity.
Figure 7: Generated images when forgetting the ‘tench’ class.
Figure 8: Ablation results with conditional DDIM on CIFAR10.
@@ Appendix B More Results
Table 7: Results on UTKFace Zhang et al. (2017) with conditional DDIM. SO: simultaneously optimizing ${\mathcal{L}}({\bm{\theta}};{\mathcal{D}}_{r})-\alpha{\mathcal{L}}({\bm{\theta}};{\mathcal{D}}_{f})$. Generated examples are shown in Fig. 10. EraseDiff achieves a better trade-off between erasing the influence of ${\mathcal{D}}_{f}$ and preserving model utility over ${\mathcal{D}}_{r}$ than SO.
Method FID over forgetting classes FID over remaining classes
$c=3\uparrow$ $c=0\downarrow$ $c=1\downarrow$ $c=2\downarrow$
Unscrubbed 8.87 7.37 11.28 9.72